complete and submit it for grading. Question 9 of 20: Select the best answer for the question 9. Construct a 95% confidence interval for p for a random sample of size n 225 yielded A. 46+.065 B. .49 t.045 C. .48 t.098 D. .42 t.082 Mark for review (Will be highlighted on the review page) Previous Question Next Question
Answer #1

Sample Size,   n =    225
      
Sample Proportion ,    p̂ =    0.46

alpha=0.05
z -value    "Zα/2 =
"   1.9600
Standard Error ,    SE = √[p̂(1-p̂)/n] =    0.0332
      
margin of error ,    E = Z*SE =    0.0651

so, confidence interval = p̂ ± E = 0.46±0.065(option a)